# Waffle: Multi-Modal Model for Automated Front-End Development

### v1

@@ Abstract
Waffle: Multi-Modal Model for Automated Front-End Development
@@ 1 Introduction
To enable MLLMs to recognize subtle differences in UI images due to minor changes in the code, we adopt contrastive learning Zhai et al. (2023); Radford et al. (2021); Gao et al. (2021) to the current task to teach MLLMs to focus on important visual differences.
@@ 2 Approach
Figure 3 represents the overview of Waffle. We create a new mutated HTML dataset (Section 2.1) for training and fine-tuning. In addition, we design structure-aware attention (Section 2.2) during model fine-tuning to teach models to focus on important HTML segments. Finally, we use contrastive learning training to teach models to learn visual differences and align the models’ visual and HMTL/CSS code understanding (Section 2.3). Specifically, we construct the training dataset by applying mutation rules for HTML code on a subset of a popular dataset, WebSight-v0.1, to generate the corresponding source code and UI images.
2.1 Training Data Mutation
To teach MLLMs the important visual differences, we create Waffle’s contrastive training data from WebSight-v0.1, which is a fine-tuning dataset built by HuggingfaceM4 and contains 822,987 pairs of HTML code and its corresponding screenshots Laurençon et al. (2024).
@@ 2.1 Training Data Mutation
We study the common mistakes of existing web MLLM, VLM-WebSight, which enabled us to create realistic mutation rules to mutate HTML/CSS code in WebSight-v0.1. We conduct a failure analysis of VLM-WebSight on 50 validation data instances. For every mismatched webpage, we manually inspect the root causes of the failures and eventually conclude seven common categories of errors as shown in Table 1. For each category, we create a set of rules to mutate an existing HTML, as shown in Appendix A.1.
Figure 4: Example of structure-aware attention.
@@ 2.2 Structure-Aware Attention
Waffle applies the structure-aware attention to the language model decoder only, and keeps the vision encoder as is. Waffle only applies the structure-aware attention mechanism to one-fourth of the attention heads in the language model decoder. This enables a quarter of attention heads to learn the structural domain knowledge explicitly, and the other three-quarters of heads to keep the standard full self-attention and pre-training knowledge. The number of attention heads that use structure-aware attention is a hyper-parameter, which can be tuned as in Section 4.2.
@@ 2.3 Contrastive-Learning
The example illustrated in Figure 2 highlights a critical gap in current models: their limited ability to effectively map variations in HTML code to the corresponding webpage screenshot. To address this, Waffle utilizes contrastive learning, allowing models to learn from comparisons and contrasts between similar examples.
More concretely, the training dataset consists of groups of HTML codes and UI Image pairs, where each group has $k$ such pairs. During training, in any group, for image $\mathbf{x}^{i}$ and code $\mathbf{y}^{i}$ ($i\in\{1,...,k\}$, image $\mathbf{x}^{i}$ is the rendered webpage image from code $\mathbf{y}^{i}$), the webpage image is split into a list of $P$$\times$ $P$ pixels patches ($P$ depends on the configuration of the backbone MLLM), each patch is encoded by the vision model $\theta_{v}$ and fused to the text model’s latent space using an adapter to result in a list of $P$ patch embeddings $\mathbf{v}^{i}=\{v^{i}_{1},v^{i}_{2},...,v^{i}_{M}\}$ ($M$ is a hyper-parameter of the backbone MLLM). The HTML code is tokenized into tokens $\mathbf{y}^{i}=\{y^{i}_{1},y^{i}_{2},...,y^{i}_{N^{i}}\}$, where $N^{i}$ is the number of the tokens of the HTML code $\mathbf{y}^{i}$. These tokens are encoded to embeddings $\mathbf{t}^{i}=\{t^{i}_{1},t^{i}_{2},...,t^{i}_{N^{i}}\}$ by the language model $\theta_{t}$ (using the structure-aware attention). We use the average over all the patch embeddings to represent the embeddings of the webpage image, and the average overall all the tokens’ embeddings to represent the embedding of the whole HTML code, denoted by:
@@
Then we calculate the cosine similarity score between the text embeddings and image embeddings, $\overline{t^{i}}$ and $\overline{v^{i}}$, bipartitely. The contrastive learning objective is to maximize the similarity between the embeddings of an HTML code and that of the corresponding UI image, which is achieved by minimizing the contrastive learning loss $L_{cl}$, along with the standard language modeling fine-tuning loss $L_{lm}$ as follows:
@@
The language modeling loss, on the other hand, aims to maximize the probability of generating the correct token $y^{i}_{j}$ given all previous token $\mathbf{y}^{i}_{<j}$ and the input webpage image $\mathbf{x}^{i}$. This is the standard objective of training an MLLM to generate text from images. Waffle jointly optimizes the two objectives as follows:
$L_{\textsc{Waffle}}=L_{lm}+\lambda L_{cl}$ (4)
where $\lambda$ is a hyper-parameter constant controlling the effect of contrastive learning on the overall optimization.
@@ 3.1 Model Training
On top of the fine-tuned MLLMs from the first step, we apply the structure-aware attention and contrastive learning approach. Structure-aware attention is applied to $\frac{1}{4}$ of the attention heads in each attention layer in the LLM decoder. Each model is trained with DoRA using the combined learning objective on the contrastive learning dataset (Equation 4 with $\lambda$ set to 0.1). The model’s weights are updated using the AdamW optimizer, with the learning rate set to $2e^{-5}$. The batch size is 32.
@@ 3.2 Test Data
Design2Code is an open-source benchmark of 484 manually processed real-world website screenshots, which are much more complex than Websight-v0.1. Evaluation on Design2Code indicates the generalization ability of models fine-tuned on Waffle’s training dataset to real-world scenarios Si et al. (2024).
@@ 3.3 Evaluation Metrics
Backbones Techniques HTML-Match (%) $\uparrow$ CW-SSIM $\uparrow$ CLIP $\uparrow$ Low-Level Element Matching (LLEM) (%) $\uparrow$
Average Block-Match Text Position Color
Gemini 1.5 Pro Prompting 9.40 0.3385 88.55 90.16 94.31 98.41 84.73 83.18
GPT-4o mini Prompting 10.20 0.3055 87.72 87.54 92.59 98.48 82.65 76.45
GPT-4o Prompting 11.40 0.3666 89.03 92.18 94.66 98.43 87.04 88.60
Moondream2 Standard FT 21.60 0.4233 89.92 90.59 91.73 96.98 87.56 86.77
Waffle 27.60 0.4486 89.98 91.72 92.26 97.25 89.55 87.81
VLM-WebSight Standard FT 28.00 0.5023 93.30 92.73 97.95 90.72 91.07 93.45
Waffle 37.00 0.6005 94.57 95.16 93.62 98.16 93.29 95.57
Table 2: Main results on the WebSight-Test dataset.
HTML-Match. HTML-Match is the percentage of generated images that match the ground truth images perfectly at the pixel level. For this comparison, styles and attributes are removed from both the ground truth and generated HTML. This process emphasizes the model’s ability to accurately recognize the text content and the DOM tree structure of the HTML code.
@@
CW-SSIM. The complex wavelet structural similarity index (CW-SSIM) computes the structural similarity between images Sampat et al. (2009).
@@
Backbones Techniques CW-SSIM $\uparrow$ CLIP $\uparrow$ Low-Level Element Matching (LLEM) (%) $\uparrow$
Average Block-Match Text Position Color
Gemini 1.5 Pro* Prompting 0.2652 87.76 87.17 91.82 97.40 82.67 76.81
GPT-4o-mini Prompting 0.2304 86.06 78.84 70.64 92.39 78.55 73.78
GPT-4o Prompting 0.2776 89.03 83.67 75.98 94.29 83.38 81.01
Moondream2 Standard FT 0.1348 46.63 40.71 29.56 49.41 40.73 43.14
Waffle 0.2142 79.62 67.83 44.32 83.59 71.61 71.81
VLM-WebSight Standard FT 0.2518 82.35 73.00 55.77 84.14 74.74 77.36
Waffle 0.2815 85.98 77.81 61.47 88.20 79.30 82.28
@@
Table 3: Main results on the Design2Code dataset.
@@ 4.1 Effectiveness of Waffle
Table 2 and Table 3 show the performance of various fine-tuning strategies on the two testing datasets, WebSight-Test (500 samples) and Design2Code (484 samples).
@@
On WebSight-Test, Waffle achieves 6.00 pp higher HTML-Match (27.60% vs. 21.60%) and 0.0253 higher CW-SSIM (0.4486 vs. 0.4233) than Standard FT with Moondream2 as the backbone. On the VLM-WebSight backbone, Waffle outperforms Standard FT by 9.00 pp HTML-Match, 0.0982 CW-SSIM, and 2.43% averaged LLEM. On Design2Code, Waffle achieves greater improvements compared to the Standard FT technique with both backbones.
Note that, Waffle is a generalizable fine-tuning pipeline that can benefit any pre-trained MLLMs.
@@
Comparison against SOTA commercial models. Due to the lack of comparable baselines, we also compare Waffle with top commercial models, which include GPT-4o mini, GPT-4o, and Gemini 1.5 Pro. We apply the direct prompting method following prior work Si et al. (2024). The result is shown in Table 2.
On the WebSight-Test dataset, models fine-tuned with Waffle perform better than the SOTA commercial models. VLM-WebSight overperforms GPT-4o by 25.60 pp on HTML-Match (37.00% vs. 11.40%) and 0.2339 on CW-SSIM (0.6005 vs. 0.3666). Similarly, for the smaller backbone, Moondream2 exceeds GPT-4o by 16.2 pp on HTML-Match (27.60% vs. 11.40%) and 0.0820 on CW-SSIM (0.4486 vs. 0.3666).
As shown in Table 3, VLM-WebSight fine-tuned by Waffle is better than GPT-4o on CW-SSIM by 0.039, and better than GPT-4o mini by 0.511 on the Design2Code dataset. However, GPT-4o is better in the other two metrics versus VLM-WebSight. Moondream2 fine-tuned by Waffle has a lower performance compared to GPT-4o and GPT-4o mini on all metrics. This is likely due to its smaller size, which could influence its generalizability to more complex, out-of-distribution data on the Design2Code dataset. Case studies are provided in Appendix A.2.
@@ 4.2 Ablation Studies
Table 4 shows the ablation study results. Waffle-${}_{\texttt{-attn}}$ brings improvements compared to the Standard FT on all metrics, and Waffle-${}_{\texttt{-contra}}$ brings a 4.40 pp improvement on HTML-Match. On the Design2Code dataset, Waffle has dominating performance on all metrics for models fine-tuned with both backbones. Across the two backbones, models fine-tuned with Waffle are higher than those fine-tuned with Waffle-${}_{\texttt{-attn}}$ by up to 0.0335 on CW-SSIM.
@@
Standard FT 07 | 17 (24) 14 | 13 (27) 17 | 18 (35) 2.90 | 2.42 (2.66)
Waffle-${}_{\texttt{-attn}}$ 15 | 16 (31) 09 | 17 (26) 24 | 16 (40) 2.55 | 2.37 (2.46)
Waffle-${}_{\texttt{-contra}}$ 38 | 20 (58) 08 | 11 (19) 10 | 15 (25) 1.67 | 2.38 (2.02)
Waffle 27 | 32 (59) 18 | 12 (30) 10 | 09 (19) 1.88 | 1.85 (1.87)
@@
Human evaluation results. We select 30 test samples from WebSight-Test and Design2Code (60 total). Each sample has four generated HTML codes and rendered webpages from Standard FT, Waffle-${}_{\texttt{-attn}}$, Waffle-${}_{\texttt{-contra}}$, and Waffle. Human raters rank the generated webpages based on similarity to the ground-truth webpage without knowing which model produced each one. Table 5 shows the human evaluation results for VLM-WebSight fine-tuned by Standard FT, Waffle-${}_{\texttt{-attn}}$, Waffle-${}_{\texttt{-contra}}$, and Waffle.
Across both datasets, Waffle has the best averaged-rankings, 1.87, outperforming both ablations and the baseline. Specifically, Waffle reaches 32 times rank 1 placement on Design2Code, showing great generalizability on more complex datasets. Waffle-${}_{\texttt{-contra}}$ is the second best technique on the two testsets, reaching 58 times rank 1, and an average ranking of 2.02. On the other hand, Waffle-${}_{\texttt{-attn}}$ is the third-best technique but still outperforms standard FT.

### v2

# Shanchao Liang

@@ Abstract
Waffle: Fine-tuning Multi-Modal Model for Automated Front-End Development thanks: To appear in the Proceedings of ACL 2025. This version is for personal use only.
@@ 1 Introduction
For the second challenge, minor variations in layout structure can still impact the content of the generated code. However, MLLMs often fail to capture these subtle differences and generate identical code for visually distinct inputs. We illustrate this issue in  Section A.2. To enable MLLMs to recognize subtle differences in UI images due to minor changes in the code, we adopt contrastive learning Zhai et al. (2023); Radford et al. (2021); Gao et al. (2021) to the current task to teach MLLMs to focus on important visual differences.
@@ 2 Approach
Figure 2 represents the overview of Waffle. We create a new mutated HTML dataset (Section 2.1) for training and fine-tuning. In addition, we design structure-aware attention (Section 2.2) during model fine-tuning to teach models to focus on important HTML segments. Finally, we use contrastive learning training to teach models to learn visual differences and align the models’ visual and HMTL/CSS code understanding (Section 2.3). We note that Waffle is a generalizable fine-tuning pipeline that can benefit any pre-trained MLLMs.
Specifically, we construct the training dataset by applying mutation rules for HTML code on a subset of a popular dataset, WebSight-v0.1, to generate the corresponding source code and UI images.
@@
2.1 Training Data Mutation
To teach MLLMs the important visual differences, we create Waffle’s contrastive training data from WebSight-v0.1, which is a fine-tuning dataset built by HuggingfaceM4 and contains 822,987 pairs of HTML code and its corresponding screenshots Laurençon et al. (2024). However, to make contrastive learning effective, the model must learn to recognize subtle yet realistic variations in UI elements.
To achieve this, we analyze common mistakes in VLM-WebSight by conducting a failure analysis on 50 validation samples, identifying seven common error categories (Table 1). Using insights from these failures, we design realistic mutation rules to modify HTML/CSS in WebSight-v0.1, ensuring that the contrastive training data reflects real-world errors. The mutation process follows the observed error frequencies; for example, since color mismatches account for 19.05% of errors (12/63), mutations on color properties are applied at the same rate (Table 1). The full set of mutation rules is detailed in Appendix A.1.
@@ 2.2 Structure-Aware Attention
Waffle applies the structure-aware attention mechanism exclusively to the language model decoder while keeping the vision encoder unchanged. Specifically, Waffle applies the structure-aware attention mask—formed by the union of parent-attention, sibling-attention, and self-attention masks—to only one-fourth of the attention heads in the decoder. This enables these heads to explicitly capture structural domain knowledge, while the remaining three-fourths retain full self-attention and leverage pre-trained knowledge. The number of attention heads using structure-aware attention is a tunable hyperparameter, which can be adjusted as described in Section 4.3.
Figure 3: Example of structure-aware attention. Yellow: self-attention. Green: parent-attention. Teal: sibling-attention. Purple: masked-out Attention.
@@ 2.3 Contrastive-Learning
To address current models’ limited ability to effectively map variations in HTML code to the corresponding UI, Waffle utilizes contrastive learning, allowing models to learn from comparisons and contrasts between similar examples.
More concretely, the training dataset consists of groups of HTML codes and UI Image pairs, where each group has $k$ such pairs. During training, in any group, for image $\mathbf{x}^{i}$ and code $\mathbf{y}^{i}$ ($i\in\{1,...,k\}$, image $\mathbf{x}^{i}$ is the rendered webpage image from code $\mathbf{y}^{i}$), the webpage image is split into a list of pixel patches. Each patch is encoded by the vision model $\theta_{v}$ and fused to the text model’s latent space using an adapter to result in a list of patch embeddings $\mathbf{v}^{i}=\{v^{i}_{1},v^{i}_{2},...,v^{i}_{M}\}$ ($M$ is a hyper-parameter of the backbone MLLM). The HTML code is tokenized into tokens $\mathbf{y}^{i}=\{y^{i}_{1},y^{i}_{2},...,y^{i}_{N^{i}}\}$, where $N^{i}$ is the number of the tokens of the HTML code $\mathbf{y}^{i}$. These tokens are encoded to embeddings $\mathbf{t}^{i}=\{t^{i}_{1},t^{i}_{2},...,t^{i}_{N^{i}}\}$ by the language model $\theta_{t}$ (using the structure-aware attention). We use the average over all the patch embeddings to represent the embeddings of the webpage image, and the average overall all the tokens’ embeddings to represent the embedding of the whole HTML code, denoted by:
@@
Then we calculate the cosine similarity score between the code and image embeddings, $\overline{t^{i}}$ and $\overline{v^{i}}$, bipartitely. The contrastive learning objective is to maximize the similarity between the embeddings of the corresponding code and UI, by minimizing the contrastive learning loss $L_{cl}$, along with the language modeling loss $L_{lm}$ as follows:
@@
The language modeling loss, on the other hand, aims to maximize the probability of generating the correct token $y^{i}_{j}$ given all previous token $\mathbf{y}^{i}_{<j}$ and the input webpage image $\mathbf{x}^{i}$. This is the standard objective of training an MLLM to generate text from images. Waffle jointly optimizes the two objectives, $L_{\textsc{Waffle}}=L_{lm}+\lambda L_{cl}$, where $\lambda$ is a hyper-parameter constant controlling the effect of contrastive learning on the overall optimization.
@@ 3.1 Model Training
On top of the fine-tuned MLLMs from the first step, we apply the structure-aware attention and contrastive learning approach. Structure-aware attention is applied to $\frac{1}{4}$ of the attention heads in each attention layer in the LLM decoder. Each model is trained with DoRA using the combined learning objective on the contrastive learning dataset ($\lambda$ set to 0.1). The model’s weights are updated using the AdamW optimizer, with the learning rate set to $2e^{-5}$. The batch size is 32.
@@ 3.2 Test Data
Design2Code is an open-source benchmark of 484 manually processed real-world website screenshots, which are more complex than Websight-v0.1. Evaluation on Design2Code indicates the generalization ability of models fine-tuned on Waffle’s training dataset to real-world scenarios.
@@ 3.3 Evaluation Metrics
We note that existing methods struggle to effectively capture subtle or structural differences in images. To address this, we adopt CLIP and Low-Level Element Matching (LLEM) metrics from previous work Si et al. (2024). Additionally, we introduce two new metrics: HTML-Match, designed to evaluate structural similarity in rendered HTML, and complex wavelet structural similarity index (CW-SSIM) to detect minor differences in images, respectively.
HTML-Match. HTML-Match is the percentage of generated images that match the ground truth images perfectly at the pixel level. For this comparison, styles and attributes are removed from both the ground truth and generated HTML. This process emphasizes the model’s ability to accurately recognize the text content and the DOM tree structure of the HTML code, and assess MLLM’s ability to recover the HTML structure from the given UI.
CW-SSIM. CW-SSIM computes the structural similarity between images Sampat et al. (2009). In certain cases, LLEM and CLIP scores might yield high scores as they failed to capture the structural difference. For example, in Figure 4 the rendered image is structurally different from the ground truth image, resulting in low CW-SSIM scores, 0.2904. However, the LLEM gives a high score of 99.74%, and the CLIP score is 90.67, which does not align with human evaluation as well as CW-SSIM.
@@
Figure 4: Different metrics on a pair of images.
In this work, we prioritize CW-SSIM as a more robust evaluation metric. Yet, we still report CLIP and LLEM as they are used in existing work.
@@
Backbones Techniques WebSight-Test Design2Code
Gemini 1.5 Pro Prompting 9.40 0.3385 88.55 90.16 0.2652 87.76 87.17
GPT-4o mini Prompting 10.20 0.3055 87.72 87.54 0.2304 86.06 78.84
GPT-4o Prompting 11.40 0.3666 89.03 92.18 89.03 83.67 75.98
Moondream2 Standard FT 21.60 0.4233 89.92 90.59 0.1348 46.63 40.71
Waffle 27.60 0.4486 89.98 91.72 0.2142 79.62 67.83
VLM-WebSight Standard FT 28.00 0.5023 93.30 92.73 0.2518 82.35 73.00
Waffle 37.00 0.6005 94.57 95.16 0.2815 85.98 77.81
@@
Table 2: Main results on the WebSight-Test and Design2Code dataset.
Figure 5: Example from WebSight-Test dataset, with the generated images by GPT-4o, Standard FT (VLM-WebSight), and Waffle (VLM-WebSight).
@@ 4.1 Effectiveness of Waffle
Table 2 show the performance of various fine-tuning strategies on the two testing datasets, WebSight-Test (500 samples) and Design2Code (484 samples).
@@
On WebSight-Test, Waffle achieves 6.00 pp higher HTML-Match (27.60% vs. 21.60%) and 0.0253 higher CW-SSIM (0.4486 vs. 0.4233) than Standard FT with Moondream2 as the backbone. On the VLM-WebSight backbone, Waffle outperforms Standard FT, i.e., original VLM-WebSight, by 9.00 pp HTML-Match, 0.0982 CW-SSIM, improving the model’s ability to generate structurally similar images and align the two modalities. On Design2Code, Waffle achieves greater improvements compared to the Standard FT technique with both backbones.
@@
Comparison against SOTA commercial models. Due to the lack of comparable baselines, we compare Waffle with top commercial models, which include GPT-4o mini, GPT-4o, and Gemini 1.5 Pro. We apply direct prompting following prior work Si et al. (2024). The result is shown in Table 2.
On the WebSight-Test dataset, models fine-tuned with Waffle perform better than the SOTA commercial models. VLM-WebSight overperforms GPT-4o by 25.60 pp on HTML-Match (37.00% vs. 11.40%) and 0.2339 on CW-SSIM (0.6005 vs. 0.3666), showing Waffle’s benefit in addressing the two challenges, i.e., generate structurally correct HTML and closing the gap between the two modalities. Similarly, for the smaller backbone, Moondream2 exceeds GPT-4o on CW-SSIM and HTML-Match.
As shown in Table 2, VLM-WebSight fine-tuned by Waffle is better than GPT-4o on CW-SSIM by 0.039, and better than GPT-4o mini by 0.511 on the Design2Code dataset. However, GPT-4o is better in the other two metrics versus VLM-WebSight. Moondream2 fine-tuned by Waffle has a lower performance compared to GPT-4o and GPT-4o mini on all metrics. This is likely due to its smaller size, which could influence its generalizability to more complex, out-of-distribution data.
@@
4.2 Case Study
Figure 5 shows the generation results for one instance from WebSight-Test. The generated webpage of GPT-4o has a CW-SSIM of 0.1353, significantly lower than that of 0.3760 from VLM-WebSight under standard fine-tuning. On the other hand, the webpage generated by VLM-WebSight fine-tuned by Waffle reaches an almost perfect CW-SSIM score. This example shows the effectiveness of using Waffle on UI-to-HTML generation.
@@ 4.3 Ablation Studies
Shown in Table 3, Waffle-${}_{\texttt{-attn}}$ brings improvements compared to the Standard FT on all metrics, and Waffle-${}_{\texttt{-contra}}$ brings a 4.40 pp improvement on HTML-Match. On the Design2Code dataset, Waffle has dominating performance on all metrics for models fine-tuned with both backbones. Across the two backbones, models fine-tuned with Waffle are higher than those fine-tuned with Waffle-${}_{\texttt{-attn}}$ by up to 0.0335 on CW-SSIM.
@@
Standard FT 09 | 34 (43) 18 | 13 (46) 34 | 25 (54) 2.88 | 2.37 (2.62)
Waffle-${}_{\texttt{-attn}}$ 22 | 23 (45) 011 | 30 (41) 22 | 26 (71) 2.67 | 2.45 (2.56)
Waffle-${}_{\texttt{-contra}}$ 60 | 33 (93) 013 | 17 (30) 11 | 26 (42) 1.78 | 2.41 (2.10)
Waffle 46 | 54 (100) 33 | 22 (55) 10 | 015 (26) 1.85 | 1.79 (1.82)
@@
Human evaluation results. We select 50 test samples from WebSight-Test and Design2Code (100 total). Each sample has four generated HTML codes and rendered webpages from Standard FT, Waffle-${}_{\texttt{-attn}}$, Waffle-${}_{\texttt{-contra}}$, and Waffle. Human raters rank the generated webpages based on similarity to the ground-truth webpage without knowing which model produced each one. Multiple results can share the same rank if they are deemed equally similar to the ground truth. Table 4 shows the human evaluation results for VLM-WebSight fine-tuned by Standard FT, Waffle-${}_{\texttt{-attn}}$, Waffle-${}_{\texttt{-contra}}$, and Waffle.
Across both datasets, Waffle has the best averaged rankings, 1.82, outperforming both ablations and the baseline. Specifically, Waffle reaches 54 times rank 1 placement on Design2Code, showing great generalizability on more complex datasets. Waffle-${}_{\texttt{-contra}}$ is the second best technique on the two testsets, reaching 93 times rank 1, and an average ranking of 2.10. On the other hand, Waffle-${}_{\texttt{-attn}}$ is the third-best technique but still outperforms standard FT.